# Evolving a Multi-Population Evolutionary-QAOA on Distributed QPUs

### v2

@@ Abstract
Our research combines an Evolutionary Algorithm (EA) with a Quantum Approximate Optimization Algorithm (QAOA) to update the ansatz parameters, in place of traditional gradient-based methods, and benchmark on the Max-Cut problem. We demonstrate that our Evolutionary-QAOA (E-QAOA) pairing performs on par or better than a COBYLA-based QAOA in terms of solution accuracy and variance, for d-3 regular graphs between 4 and 26 nodes, using both max_count and Conditional Value at Risk (CVaR) for fitness function evaluations. Furthermore, we take our algorithm one step further and present a novel approach by presenting a multi-population EA distributed on two QPUs, which evolves independent and isolated populations in parallel, classically communicating elite individuals. Experiments were conducted on both simulators and IBM quantum hardware, and we investigated the relative performance accuracy and variance.
@@ I Introduction
Quantum Approximate Optimization Algorithms (QAOAs) are promising methods for solving combinatorial optimization problems [1, 2, 3] on quantum computers. The original QAOA presented by Farhi, Goldstone and Gutmann in 2014 [4] demonstrated the potential for quantum advantage as it outperformed the best-known classical algorithm to solve the Max-Cut problem at the time. Since then classical computing has taken greater strides to maintain an advantage over quantum optimization. Yet, the QAOA continues to be an active area of research for the quantum computing community [5]. The QAOA is a variational algorithm [6] and therefore faces the well-known challenges of barren plateaus and noise degradation [7, 8] that occur with increasing circuit width and depth in the current era of noisy, near-term quantum computers. These issues make it increasingly hard for optimizers to move in a search landscape as the problem grows to a relevant scale [9, 10]. These difficulties motivate our research into an Evolutionary Algorithm (EA) as an optimizer for QAOA, instead of linear or gradient-based methods. There is still a discussion within the literature about whether or not gradient-free methods are useful, for example, Arrasmith et al. argue that these do not solve the barren plateau problem [11]. However, others state that Genetic Algorithms (GAs) and EAs have an advantage for faster convergence [12], avoiding barren plateaus [13], and beating the state-of-the-art gradient-free optimization for better accuracy [14]. Note that there is no clear consensus in the field of the distinction between GA and EA and the two have overlapping definitions [15]. As such, a benefit for one would clearly benefit the other. Furthermore, EAs have been shown to handle complex optimization problems with noisy or uncertain fitness functions [15, 16]. If these conditions hold, and EAs prove to be more robust as the literature continues to tackle larger-scale problems, this would make them a natural choice in combining them with QAOAs and other Variational Quantum Algorithms (VQAs).
@@
In addition, EAs offer the potential for embarrassingly parallel applications, one specific method developed is that of Multi-population Genetic Algorithms (MGAs) [17, 18], which evolve independent populations in isolation. MGAs have shown to be successful in avoiding premature convergence, an issue that arises sometimes in GAs, and increasing diversity within the population’s evolution, a key factor for successful convergence [17, 19, 20]. We present a technique that implements an MGA, combining a multi-population evolutionary algorithm with a QAOA. We propose to distribute independent population instances on two separate Quantum Processing Units (QPUs) orchestrated by a scheduler. This strategy is novel to the best of the authors’ knowledge: numerous scheduling prescriptions have been proposed by the community [21, 22, 23], however, none are tailored to parallelizing MGAs across QPUs.
@@ II Methodology
In this study, we present the components of our hybrid evolutionary QAOA, referred to as E-QAOA. Firstly, we describe the problem we are trying to solve, the Max-Cut problem. Secondly, we detail how to construct the quantum circuit and how to process measurement outcomes. Thirdly, we give an overview of the EA elements required to evolve a population of solutions. Note that this section will also give a brief overview of the terminology used in the fields of evolutionary and genetic algorithms, to then use them throughout the rest of the paper. It is worth mentioning that in the EA literature, the functional cost is also referred to as fitness; as such these two terms will be used interchangeably.
@@ II-A Max-Cut
As mentioned already, the first QAOA published in the literature [4] used a variational quantum circuit to solve the Max-Cut problem. Where the solution to this problem can be neatly summarized as stating that given an unweighted graph, with $n$ nodes, and $e$ edges, the Max-Cut of this graph is given by the boundary which divides the nodes into two sets, such that the number of edges between the two sets is maximized. In the quantum setting, there is a one-to-one mapping between nodes and qubits. The quantum algorithm samples bit-strings $x\in\{0,1\}^{n}$ where the set that a given node belongs to is labelled as ‘0’ or ‘1’.
@@ II-B The QAOA circuit
Figure 1: Example circuits for $d$-3 regular graphs with 4 (top) and 6 (bottom) nodes. The circuits have as many qubits as there are in the respective graphs, and are structured as follows: a Hadamard gate ($H$, in red) is first applied to all qubits, followed by $RZZ$ gates constituting the connectivity section and $Rx$ gates on all qubits. This circuit section (dotted-line rectangle) is repeated $p$ times, before the qubit states are measured. Crucially, increasing either $p$ or the number of edges, i.e. the graph connectivity, increases the circuit depth.
We chose a basic QAOA without enhancements, to effectively study the effects of pairing an EA to it. The version we chose is described in Qiskit’s documentation [27]. In the plain algorithm, the circuit is composed of parameterised time evolution unitaries that are alternately generated by a cost Hamiltonian ($H_{C}$) and a mixer Hamiltonian ($H_{M}$). Given the Max-Cut problem, as described above, the unitaries of these are given by
@@
where $X_{i}$ and $Z_{i}$ are the Pauli X and Z matrices respectively, acting on the $i^{\rm th}$ qubit (node). The $RZZ$ gates are applied to qubit pairs $(i,j)$ that correspond to edges in the graph. The angle of these rotations is determined by the ansatz parameters $\beta$ and $\gamma$. The unitary $U(H_{M})$ and $U(H_{C})$ blocks are repeated $p$ times, where $p$ is also known as the number of rounds in QAOA. As $\lim_{p\to\infty}$, the algorithm converges to the optimal solution, given the correct optimization of the ansatz parameters. The parameters $\beta$ and $\gamma$ change for each $p$ block, and are initialized randomly on $(-\pi,\pi]$. The resulting circuit structure for a given graph is aided by a simple diagram in Fig. 1. While many optimizers can be chosen for QAOA, the traditional choice in the literature and documentation used to update $\beta$ and $\gamma$ is COBYLA, which is a linear-based optimiser, and in this work, it was used through the minimize method in Scipy [28].
@@ II-C Fitness evaluation
In our experiments, we considered two metrics to measure the fitness of the solution, and thus update $\beta$ and $\gamma$. In the first instance, the fitness was defined as the Max-Cut number of the most frequent bit-string, rather than the average or expected output, given by the measured distribution of each circuit simulation. This can be obtained through Qiskit’s max_count feature. In the second instance, we used Conditional Value-at-Risk (CVaR), which takes the expectation value of the best $\alpha$ partitions found in the output distribution [29]. Where $\alpha$ is a value between 0 and 1, and $\alpha=1$ is equivalent to the standard expectation value using the whole distribution. We chose this method as CVaR leads to faster convergence on both simulators and hardware, as tested on various combinatorial optimization problems [30]. Finally, rather than using the Max-Cut value of these partitions directly, we use the approximation ratio, which is simply the Max-Cut value we find divided by the true optimal Max-Cut solution, such that this value is normalized on 1 for all graph sizes, and we can see how close we are to the true solution:
[…]
II-D The Evolutionary algorithm
In our method, we replace COBYLA with an EA that is used to update the variational parameters. All evolutionary or genetic algorithms have some core components to their structure and may slightly change in things like mutation operations or reproduction techniques, but the field allows the two to be used interchangeably. The core components describing an EA are thus: parent selection, recombination operator, mutation, and survivor selection. These choices of strategy help maintain uniqueness and diversity within the population genome such that the algorithm can achieve a good balance between exploitation and exploration in its evolution and research of the fitness landscape. There are other factors as well that can compose an evolutionary algorithm, like population size, replacement strategy, termination strategy and more, but these are less important than the core components. We will limit our explanation of the background theory of this field to the components that we made use of, and we will lead the readers through the terminology that we mention. In an EA we start with a population of solutions, so rather than just one initial random point for most optimization methods, we start with $n_{\rm pop}$ values. These are also initialized randomly, and each individual in the population, denoted by $Y=\{Y_{1},Y_{2},\cdots,Y_{n_{\rm pop}}\}$, is described the same as in COBYLA, with a real-valued array between $-\pi$ and $\pi$ of the form
$Y_{i}=[\beta_{1},\gamma_{1},\beta_{2},\gamma_{2},\cdots,\beta_{p},\gamma_{p}]\,.$ (4)
In evolutionary terminology, this array is defined as the genotype or chromosome, and all of the values within this array are referred to as genes or alleles. As the reader will note, the terminology within the field adopts terms from biological evolution theory. In the metaphor that compares Darwinian evolution to a programmed evolution, GAs use these notions which make it easier for the reader and users of it to understand the concepts being used, how they form the algorithm, and why they should operate. Thus, the value of $Y$ is described as the genotype because in evolution the genetic code of an individual will be the factor determining its fitness relative to the population, its likelihood to survive and reproduce and carry forth its genetic code in future generations. From this, the algorithm follows. Once we have our population, we determine the fitness of each $Y_{i}$ and use that value to choose the individuals that will pair up as parents and reproduce to form a new generation of solutions or offspring. The parent selection strategy is given by Stochastic Universal Sampling (SUS) [31] where each individual’s probability of being chosen as a parent is proportional to their normalized fitness value. This ensures that all individuals have a chance to carry their genes over to the next generation, apart from the weakest link. We have chosen for the population size to remain constant, such that two parents produce two new offspring until a population of the same size is achieved. SUS allows sampling with replacement, where the fittest individual can be chosen in multiple, or all pairs of parents, but can only be selected once for each pair of parents and thus they cannot pair with themselves. Once the parents are chosen, the offspring are produced via recombination and mutation. Recombination was applied $100\%$ of the time, on each pair of parents, through whole arithmetic crossover, i.e. on all of their genes. This operation is described below, where, given two parents, $Y_{i}$ and $Y_{j}$, one of their offspring $Y_{i}^{\prime}$ is given by
@@ II-D The Evolutionary algorithm
where $\alpha$ is a randomly generated variable uniform on $(0,1]$. Mutation is then applied to each of the genes with a probability $p_{\sigma}=0.2$, this again follows the concept of natural evolution, where offspring are never the exact combination of their parents, but in fact, undergo random mutation on their genetic codes. In our EA we use self-adaptive mutation, where we add $2p$ different $\sigma$ mutation values within the genotype, such that they also undergo recombination and mutation through the generations, and they evolve alongside the weights $\beta$ and $\gamma$. By attaching these $\sigma$ values, the genotypes become
@@
Each $\sigma$ is normally distributed on $(0,1]$ and mutation occurs as follows:
@@
The variables $\tau$ and $\tau^{\prime}$ are usually set to be inversely proportional to the square root of the population size but can be defined by the user, and can be seen as a learning rate [15]. Furthermore, $\sigma$ is mutated by a log-normal distribution to meet certain requirements, like smaller modifications, the equal likelihood of drawing a value and its reciprocal [32]. After these values evolve we enforce a minimum threshold such that $|\sigma^{\prime}|\geq\sigma_{\rm min}$. If recombination or mutation parameters push $\beta$ and $\gamma$ out of the predefined ranges, we enforce periodic boundaries to return the genes to their proper domain.
After the new generation is created we implement generational survivor selection, where $\mu$ fittest individuals from the older generation will be preserved in the newer generation, if no offspring can surpass their fitness values. This ensures that the best-performing individuals are not lost in evolution and that their genes continue to be passed on through generations. This strategy is particularly useful in noisy environments like quantum hardware, where this safeguards a standard fitness, in such a case even if the population’s average fitness happens to decrease, the next generations can always recuperate and spread the fittest genes again. The operations described are repeated in this order for a fixed number of generations, $g$. Finally, the pseudo-code for our E-QAOA is summarized in the steps of Algorithm 1.
Algorithm 1 : The steps for evolving a population
@@
The approach for a multi-population EA then follows. Each QPU evolves an independent population of solutions, as we have described above, for a fraction of the total number of generations $g_{\rm f}$, e.g. $g_{\rm f}=g/3$, at which point the fittest individuals from each population migrate to their non-native population. In our case, the fittest individuals are copied over, and they replace the weakest individuals to maintain a constant population size. In this case, the fittest individuals to migrate are chosen by looking at the absolute best Max-Cut given by their distribution, rather than their $\alpha$-tail fitness given by CVaR. The weakest individuals are also evaluated this way before being replaced. Once these switches are done, the independent evolutions continue as normal, with migrations happening every $g_{\rm f}$ generations until termination. These migrations happen classically and their execution is therefore affected by potential queue times and scheduling distributions manually. As mentioned earlier, one of the properties of an MGA approach is to increase diversity within a population. We measure this diversity by looking at uniqueness of values, in particular the number of unique fitness values and unique genes across the population. We normalize this value by the size of the population:
Uniqueness ratio=Number of unique values $\textrm{Uniqueness~{}ratio}=\frac{\text{Number of unique values found}}{n_{\rm pop}}.$ (10)

### v3

@@ Abstract
Our work integrates an Evolutionary Algorithm (EA) with the Quantum Approximate Optimization Algorithm (QAOA) to optimize ansatz parameters in place of traditional gradient-based methods. We benchmark this Evolutionary-QAOA (E-QAOA) approach on the Max-Cut problem for d-3 regular graphs of 4 to 26 nodes, demonstrating equal or higher accuracy and reduced variance compared to COBYLA-based QAOA, especially when using Conditional Value at Risk (CVaR) for fitness evaluations. Additionally, we propose a novel distributed multi-population EA strategy, executing parallel, independent populations on two quantum processing units (QPUs) with classical communication of ‘elite’ solutions. Experiments on quantum simulators and IBM hardware validate the approach. We also discuss potential extensions of our method and outline promising future directions in scalable, distributed quantum optimization on hybrid quantum-classical infrastructures.
@@ I Introduction
Quantum Approximate Optimization Algorithms (QAOAs) [1] are promising methods for solving combinatorial optimization problems [2, 3, 4, 5, 6] on quantum computers. Some of the first implementations of QAOA [7, 8] demonstrated the potential for quantum advantage as it outperformed the best-known classical algorithm to solve the Max-Cut problem at the time, assuming the use of optimal parameters [9]. Since then, classical computing has taken greater strides to maintain an advantage over quantum optimization. Yet, the QAOA continues to be an active area of research for the quantum computing community [10]. The QAOA is a variational algorithm [11, 12] and therefore faces the well-known challenges of barren plateaus and noise degradation [13, 14] that occur with increasing circuit width and depth in the current era of noisy, near-term quantum computers. These issues make it increasingly hard for optimizers to move in a search landscape as the problem grows to a relevant scale [15, 16]. These difficulties motivate our research into an Evolutionary Algorithm (EA) as an optimizer for QAOA, instead of linear or gradient-based methods. There is still a discussion within the literature about whether or not gradient-free methods are useful, for example, Arrasmith et al. argue that these do not solve the barren plateau problem [17]. However, others state that Genetic Algorithms (GAs) and EAs have an advantage for faster convergence [18], avoiding barren plateaus [19], and beating the state-of-the-art gradient-free optimization for better accuracy [20]. Note that there is no clear consensus in the field on the distinction between GA and EA and the two have overlapping definitions [21]. As such, a benefit for one would clearly benefit the other. Furthermore, EAs have been shown to handle complex optimization problems with noisy or uncertain fitness functions [21, 22]. If these conditions hold, and EAs prove to be more robust as the literature continues to tackle larger-scale problems, this would make them a natural choice in combining them with QAOAs and other Variational Quantum Algorithms (VQAs).
@@
In addition, EAs offer the potential for embarrassingly parallel applications, one specific method developed is that of Multi-population Genetic Algorithms (MGAs) [23, 24], which evolve independent populations in isolation. MGAs have been shown to be successful in avoiding premature convergence, an issue that sometimes arises in GAs, and increasing diversity within the population’s evolution, a key factor for successful convergence [23, 25, 26]. We present a technique that implements an MGA, combining a multi-population evolutionary algorithm with a QAOA. We propose to distribute independent population instances on two separate Quantum Processing Units (QPUs) orchestrated by a scheduler. This strategy is novel to the best of the authors’ knowledge: numerous scheduling prescriptions have been proposed by the community [27, 28, 29]; however, none are tailored to parallelizing MGAs across QPUs.
@@ II Methodology
In this study, we propose a new hybrid framework, termed the Evolutionary Quantum Approximate Optimization Algorithm (E-QAOA), which combines variational quantum algorithms and classical evolutionary optimization. The methodology is organized into three main components, described as follows.
  1. 1.
    Formulation of the Max-Cut problem, a well-known combinatorial optimization task.
  2. 2.
    Design of the QAOA circuit, consisting of state preparation, unitary evolution, and post-processing of measurement outcomes.
  3. 3.
    Evolutionary strategies used to optimize the variational parameters, discussing key concepts and terminologies from evolutionary and genetic algorithms.
Notably, in the evolutionary computation literature, the terms cost and fitness are used interchangeably; both refer to the objective function guiding the optimization.
@@ II-A Max-Cut
As demonstrated in the seminal work by Farhi et al. [1], QAOA can be adapted to solve the Max-Cut problem by optimizing a variational quantum circuit. For an unweighted graph $G=(V,E)$ with $n$ nodes, the goal is to partition $V$ into two subsets that maximize the number of edges between them. Equivalently, the objective is to maximize $C(x)=1/2\,\sum_{(i,j)\in E}1-(1-2x_{i})(1-2x_{j})$, where each $x_{i}\in\{0,1\}$ indicates the partition assignment of node $i$. In the quantum setting, nodes are mapped one-to-one to qubits, and the algorithm samples bit-strings $x\in\{0,1\}^{n}$, with each bit representing the corresponding node’s partition.
@@ II-B The QAOA circuit
Figure 1: Example circuits for $d$-3 regular graphs with 4 (top) and 6 (bottom) nodes. The circuits have as many qubits as there are in the respective graphs, and are structured as follows: a Hadamard gate ($H$, in red) is first applied to all qubits, followed by $RZZ$ gates constituting the connectivity section and $Rx$ gates on all qubits. This circuit section (dotted-line rectangle) is repeated $p$ times before the qubit states are measured. Crucially, increasing either $p$ or the number of edges, i.e. the graph connectivity, increases the circuit depth.
We adopt the standard QAOA, as implemented in Qiskit [33], to analyze the performance improvement when combined with an EA. In our implementation, the circuit alternates between two parameterized unitary operators. The mixer and cost unitaries are given by
@@
respectively, where $X_{i}$ and $Z_{i}$ denote the Pauli-$X$ and -$Z$ operators on the $i^{\rm th}$ qubit, and the corresponding $RZZ$ gates are applied to each edge $(i,j)$ in the graph. These layers are repeated $p$ times (with each layer having distinct parameters $\beta$ and $\gamma$, initialized randomly on $(-\pi,\pi]$). A schematic of the circuit for a given graph is presented in Fig. 1. In the limit $p\to\infty$, QAOA converges to the optimal solution and finds the global cost minimum under ideal parameter optimization.
Finally, we use COBYLA (via the minimize method in Scipy [34]) to update $\beta$ and $\gamma$ as our control experiment. COBYLA has been shown to provide generally good convergence, on par with other common optimizers like SLSQP, L‑BFGS‑B [35]. For this reason, we restrict our investigation to COBYLA, with a plan to compare EAs to other optimizer choices (e.g., as in [36]) in future work.
@@ II-C Fitness evaluation
In our experiments, we employ two metrics to evaluate the fitness of candidate solutions and update the QAOA parameters $\beta$ and $\gamma$. First, fitness is defined as the Max-Cut value corresponding to the most frequent bit-string from the measurement distribution, obtained via Qiskit’s max_count feature. Second, we use Conditional Value-at-Risk (CVaR) [37], which computes the expectation value over the best $\alpha$ fraction of outcomes (with $\alpha\in(0,1]$, where $\alpha=1$ yields the standard expectation). Empirical studies have shown that CVaR leads to faster convergence on both simulators and quantum hardware [38] with respect to using the standard expectation value, as the optimization is driven by the expectation of the $\alpha$-tail of the distribution rather than the whole distribution, or a single bit-string (max_count). In the latter method, the number of shots required to reliably converge using the most frequent occurring bit-string would scale combinatorially with the number of nodes for our problem, quickly saturating the capabilities of near-term quantum hardware and ultimately returning a randomly selected bit-string due to noise. For this reason, we only show results using the max_count bit-string selection in Fig. 2 to illustrate the comparison with CVaR, and use the CVaR method throughout the rest of the paper.
To compare performance across various graph sizes, we define the approximation ratio, $\mathcal{R}$, as
$\mathcal{R}=\frac{C({\rm solution~{}found})}{C({\rm optimal~{}solution})},$ (3)
which equals 1 when the optimal solution is achieved.
II-D Evolutionary algorithm
Our EA implementation features the core rules of evolutionary dynamics: parent selection, recombination operator, mutation, and survivor selection. These components ensure a balance between exploration of the cost (or fitness) landscape and exploitation around candidate solutions. Additional heuristic rules and properties can be added to the EA to improve the convergence performance (e.g. replacement and termination strategy), however, these are less crucial than the core rules and will not be investigated in this work.
The process begins with a population of $n_{\rm pop}$ candidate solutions, denoted by $Y=\{Y_{1},Y_{2},\cdots,Y_{n_{\rm pop}}\}$, each encoded as a ‘genotype’ (or chromosome) consisting of real-valued parameters:
$Y_{i}=[\beta_{1},\gamma_{1},\beta_{2},\gamma_{2},\ldots,\beta_{p},\gamma_{p}],$ (4)
with each gene (or allele) initialized randomly in $[-\pi,\pi]$.
Given these initial conditions, the algorithm updates the population by selecting ‘parent’ solutions to produce ‘offsprings’. Parents are selected using Stochastic Universal Sampling (SUS) [39], where the selection probability is proportional to the normalized fitness values. This ensures that all individuals $Y_{i}$ have a chance to maintain their genes through the following generation, except for the least fit solution. We keep the population size constant by imposing that two parents generate two new offspring. SUS allows sampling with replacement, where the fittest individual can be chosen in multiple (or all) pairs of parents but can only be selected once for each pair of parents, and thus they cannot pair with themselves. Once the parents are chosen, the offspring are produced via recombination and mutation.
Recombination was applied $100\%$ of the time on each pair of parents through whole arithmetic crossover, i.e. on all of their genes. This operation is described as follows. Given two parents, $Y_{i}$ and $Y_{j}$, one of their offspring $Y_{i}^{\prime}$ is given by
@@ II-D Evolutionary algorithm
where $\alpha$ is a randomly generated variable uniform on $[0,1]$ (not to be confused with the CVaR $\alpha$ parameter).
Mutation is then applied to each of the genes with a probability $p_{\sigma}=0.2$. This heuristic strategy is, once again, inspired by Darwinian natural evolution, where offspring are most likely never the exact combination of their parents but are subject to random mutation on their genome. In our EA, we use self-adaptive mutation, where we add $2p$ different $\sigma$ mutation values within the genotype, such that they also undergo recombination and mutation through the generations, and they evolve alongside the weights $\beta$ and $\gamma$. By attaching these $\sigma$ values, the genotypes become
@@
Each $\sigma$ is normally distributed on $[0,1]$ and mutation occurs as follows:
@@
The variables $\tau$ and $\tau^{\prime}$ are usually set to be inversely proportional to the square and fourth root of the population size, respectively, but can be defined by the user, and can be interpreted as a learning rate [21]. Furthermore, $\sigma$ is mutated by a log-normal distribution to meet certain requirements, like smaller modifications, the equal likelihood of drawing a value and its reciprocal [40]. After these values evolve we enforce a minimum threshold such that $|\sigma^{\prime}|\geq\sigma_{\rm min}$. If recombination or mutation parameters push $\beta$ and $\gamma$ out of the predefined ranges, we enforce periodic boundaries to return the genes to their proper domain.
After creating the new generation, we implement generational survivor selection, where $\mu$ fittest individuals from the older generation will be preserved in the newer generation if no offspring can surpass their fitness values. This ensures that the best-performing individuals are not lost in evolution and that their genes continue to be passed on through generations. This strategy is particularly useful in the presence of noise, like currently available quantum hardware: even if the population’s average fitness decreases, the next generations can retain and spread the fittest genes again. The operations described are repeated in this order for a fixed number of generations, $g$. Finally, the pseudo-code for our E-QAOA is summarized in Algorithm 1.
Algorithm 1 Steps to implement this work’s EA for evolving a population of solutions.
@@
The approach for a multi-population EA is similar. In our set-up, we consider multiple QPUs, each evolving an independent population of solutions, as described above, for a fraction $g_{\rm f}$ of the total number of generations, e.g. $g_{\rm f}=g/3$. At every $g_{\rm f}$ iterations, the fittest individuals from each population migrate to their non-native population. In our computational set-up, the fittest individuals are copied to the target device to replace the weakest individuals and maintain a constant population size. At migration time, the fittest individuals to migrate are selected based on the absolute best Max-Cut value in the distribution rather than their $\alpha$-tail fitness given by CVaR. The weakest individuals are also evaluated this way before being replaced. Then, the populations continue their evolution independently until the next migration or upon reaching the maximum number of iterations required for termination.
When implemented on a real quantum-classical infrastructure, migration occurs classically so that the job is re-queued. Therefore, the overall execution time is impacted by the overall system usage, network traffic and, where necessary, manual scheduling of restart jobs.
As mentioned above, one of the benefits of an MGA approach is an increased diversity within a population of solutions, boosting the chances of guiding the algorithm towards optimality. We measure diversity by quantifying the uniqueness of values, in particular the number of unique fitness values and unique genes across the population. We normalize this value by the size of the population as follows:
Uniqueness $\textrm{Uniqueness~{}ratio}=\frac{|\{\beta(Y_{i}),\gamma(Y_{i}):Y_{i}\in Y\}|}{n_{\rm pop}},$ (10)
where ||$\cdot$|| ⋅ | denotes the umber of unique gene values found in the population.